OpenStudy (anonymous):

Let calculate the values of integral. For α>0.

OpenStudy (anonymous):

Let calculate the values of integral. For \(\alpha > 0\) \[\int\limits_{0}^{\infty} -x e^{- \alpha x} dx\]

OpenStudy (anonymous):

use product rule : (uv)'=u'v+uv' let u=-x and dv=e^{-ax}dx make this substitution and use the above product rule

OpenStudy (anonymous):

is this for Laplace transform? :)

OpenStudy (anonymous):

how about this \[-\frac{ 1 }{ \alpha^2 } \left( - \alpha x - 1 \right) e^{- \alpha x} |_{0}^{\infty}\]

OpenStudy (anonymous):

no isn't :)

OpenStudy (anonymous):

@myko

OpenStudy (anonymous):

hbu @RadEn ??

OpenStudy (raden):

use the table integral, the result of that integration is